# Perspective on a challenge: predicting the photochemistry of cyclobutanone

### 14 Apr 2026 v1

@@ 4 Summary of the results and Outlook
Finally, coming back to the original question that motivated this challenge – Is the field of nonadiabatic molecular dynamics mature enough to be predictive? – one may be tempted to answer nearly. This exercise showed that nonadiabatic molecular dynamics can be qualitatively predictive: Fig. 7 showed that many predicted UED signals reproduced the main features of the experimental UED time trace and would provide valuable insight to guide the analysis of the experimental signal. Nearly quantitative reproduction of major experimental UED features was achieved with nonadiabatic dynamics using multireference electronic-structure methods. From a chemical perspective, most contributions using a multiconfigurational/multireference approach obtained the same families of photoproducts, mostly formed on the ground electronic state following nonradiative decay. All contributions agreed that intersystem crossing processes are not important for the photochemistry of gas-phase cyclobutanone photoexcited at 200 nm. The discussions contained in this Perspective also provide important guidelines for benchmarking that should support the predictive power of future nonadiabatic dynamics simulations of gas-phase molecules (such guidelines are also further detailed in recent worksPrlj et al. (2026)). Nevertheless, we contrast this positive outcome with more realistic considerations. First, this challenge focused on a single experimental observable from MeV-UED measurements, which may have been more forgiving to some methods than other experimental observables would have been. Second, this challenge encompassed both photophysical and photochemical processes, which constitute a stringent test for nonadiabatic molecular dynamics. However, these processes took place in the gas phase –advances in liquid jets at advanced light sources mean that the nonadiabatic dynamics community may soon need to consider similar tests for ultrafast dynamics in complex environments.

### 16 Apr 2026 v2

@@ 4 Summary of the results and Outlook
Finally, coming back to the original question that motivated this challenge – Is the field of nonadiabatic molecular dynamics mature enough to be predictive? – one may be tempted to answer nearly. This exercise showed that nonadiabatic molecular dynamics can be qualitatively predictive: Fig. 7 showed that many predicted UED signals reproduced the main features of the experimental UED time trace and would provide valuable insight to guide the analysis of the experimental signal. Nearly quantitative reproduction of major experimental UED features was achieved with nonadiabatic dynamics using electronic-structure methods including both static and dynamic correlation. From a chemical perspective, most contributions using an electronic-structure approach including static correlation obtained the same families of photoproducts, mostly formed on the ground electronic state following nonradiative decay. All contributions agreed that intersystem crossing processes are not important for the photochemistry of gas-phase cyclobutanone photoexcited at 200 nm. The discussions contained in this Perspective also provide important guidelines for benchmarking that should support the predictive power of future nonadiabatic dynamics simulations of gas-phase molecules (such guidelines are also further detailed in recent worksPrlj et al. (2026)). Nevertheless, we contrast this positive outcome with more realistic considerations. First, this challenge focused on a single experimental observable from MeV-UED measurements, which may have been more forgiving to some methods than other experimental observables would have been. Second, this challenge encompassed both photophysical and photochemical processes, which constitute a stringent test for nonadiabatic molecular dynamics. However, these processes took place in the gas phase –advances in liquid jets at advanced light sources mean that the nonadiabatic dynamics community may soon need to consider similar tests for ultrafast dynamics in complex environments.